a source is moving toward observer with velocity c/2 and observer is also moving in same direction with velocity c/3 if the actual frequency of light emmited by source is v0 what is the apperant frequency of light measured by observer
1
Expert's answer
2020-09-08T09:11:09-0400

"\nu = \nu_0 \frac{\sqrt{1-\frac{v^2}{c^2}}}{1-\frac{v}{c}cos\theta}"

where "v" is a velocity of source relative to the observer, "\theta" is the angle between directions of source and observer. When source is behind the observer, "\theta=0"; when the source will be ahead, "\theta=\pi".

To find relative velocity, we use relativistic velocity-addition formula:

"v = \frac{u_2-u_1}{1-\frac{u_1u_2}{c^2}}= \frac{c/2-c/3}{1-1/6}= \frac{c}{5}"

Before the moment of meeting (source is coming closer):

"\nu = \nu_0 \frac{\sqrt{1-\frac{1}{5^2}}}{1-\frac{1}{5}}= \nu_0 \frac{\sqrt{24}}{5}\frac{5}{4} = 1.225 \nu_0"

After the moment of meeting (source becomes more and more distant from the observer):

"\nu = \nu_0 \frac{\sqrt{1-\frac{1}{5^2}}}{1+\frac{1}{5}}= \nu_0 \frac{\sqrt{24}}{5}\frac{5}{6} = 0.816\nu_0"